# PPFL-RDSN: Privacy-Preserving Federated Learning-based Residual Dense Spatial Networks for Encrypted Lossy Image Reconstruction

### v1

@@ 3.1 Threat Modeling and Mitigation
  • Local Differential Privacy: Our carefully tailored additive DP noise is applied to the local model training process so that this can prevent the Inference Attacks mentioned above. The detailed design of LDP is described in Section 3.2.
  • Model Watermarking Mechanism: As described in Section 3.3, the proposed watermarking mechanism prevents model substitution attacks or unauthorized alterations by the aggregator. It detects attempts at model substitution, enabling clients to verify the integrity of global model updates and ensuring individual contributions remain uncompromised.
3.2 Local Differential Privacy (LDP)
@@
Step 1: Data‑independent orthogonal transform.
For a fixed layer $k$ we denote its activation:
$H=g_{k}(\dots,X;W_{k})\in\mathbb{R}^{C\times h\times w}.$
where $g_{k}$ uses its parameters $W_{k}$ to transform the output of the previous layer, producing an activation tensor $H$ with dimensions $C\times h\times w$. We apply a data‑independent orthogonal transform $C=THT^{\!\top},$ where $T$ is a 2‑D Discrete Cosine Transform (DCT). Orthogonality ensures $\|C\|_{2}=\|H\|_{2}$, which simplifies sensitivity analysis.
@@
Setting $\tau$ before training means the partition is input‑agnostic and does not consume any privacy budget.
The coefficients in $\mathcal{S}_{L}$ (global structure) are retained, whereas those in $\mathcal{S}_{H}$ (fine detail) will be noised. Writing $C_{L}$ and $C_{H}$ for the masked tensors and transforming back, $H_{L}=T^{\!\top}C_{L}T$, $H_{H}=T^{\!\top}C_{H}T$.
@@
By the Gaussian mechanism [54] the released $\widetilde{f}$ is $(\epsilon,\delta)$‑DP. The noisy activation is reconstructed as
@@
Algorithm.
The full client routine is summarized in Algorithm 1. Multiple rounds are composed via the moments accountant; in our experiments $30$ local steps with $\sigma{=}0.8$ yield a total $\epsilon\!\approx\!3$ at $\delta=10^{-5}$.
@@
2:$H\leftarrow g_{k}(\dots,X;W_{k})$triangleright Forward to layer $k$
3:$\widehat{H}\leftarrow\textsc{ClipL2}(H,C_{\max})$
4:$C\leftarrow T\widehat{H}T^{\!\top}$triangleright Fixed orthogonal transform
5:$(C_{L},C_{H})\leftarrow\textsc{Partition}(C,\tau)$
6:$f\leftarrow\operatorname{vec}(C_{H})$
7:$\sigma\leftarrow(2C_{\max}\sqrt{2\ln(1.25/\delta)})/\epsilon$
8:$\widetilde{f}\leftarrow f+\mathcal{N}(0,\sigma^{2}I)$
@@
In summary, our client‑side LDP module safeguards the input data while retaining reconstruction quality: we first partition DCT‑domain activations with a publicly fixed threshold, incurring no privacy cost or an explicitly budgeted one if adaptively learned; we then clip the activations and exploit the norm‑preserving orthogonality of the DCT to derive a tight global $L_{2}$ sensitivity of $2C_{\max}$; finally, Gaussian noise calibrated to this bound delivers formal $(\epsilon,\delta)$‑DP guarantees, with privacy loss composed across rounds via the moments accountant. Because noise is applied exclusively to high‑frequency coefficients, the low‑frequency content that dominates perceptual fidelity remains intact, thereby achieving strong privacy protection and high‑quality image reconstruction within the PPFL‑RDSN framework.
3.3 Model Watermarking
To ensure provenance and integrity of local updates in the PPFL–RDSN framework, we embed a client‑specific digital watermark that (i) does not degrade task performance, (ii) survives federated aggregation, and (iii) remains detectable even in very high‑dimensional models. Our design draws on two strands of priors: parameter‑regularizer watermarking [55] and bi‑level optimisation for robustness [56]. Empirically, such schemes incur $<0.05\%$ accuracy loss while withstanding pruning of $65\%$ weights or 10 epochs of fine‑tuning.
@@ 3.3 Model Watermarking
where $d$ is the model dimension and $p\!\in\!(10^{-4}-10^{-2})$. We reuse the same secret key $mk$, which makes every client perturb an identical, extremely sparse coordinate subset $J$ chosen by a PRF, so their updates reinforce each other and the watermark emerges clearly only in the aggregated model. This coordination is secure because the PRF hides $J$, making the touched indices computationally indistinguishable from random; the sparsity parameter $p\leq 10^{-2}$ ensures each client’s watermark is lost in ordinary gradient noise; and the perturbations remain negligible in any single update, so even a large ground of attackers would struggle to uncover $J$. Thus, sharing the same coordinates is a deliberate design that maximizes watermark detectability while preserving secrecy and robustness.
@@
Instead of post‑hoc noise injection, we jointly optimise the task loss $\mathcal{L}_{\text{task}}$ and a watermark loss
$\mathcal{L}_{\text{wm}}=\frac{1}{|J|}\sum_{j\in J}(\theta_{j}w_{j}-\beta)^{2},$
$\text{with }w_{j}\!=\!\text{PRF}(mk,(w,j)),\;\beta>0,$
which drives the correlation between $\theta_{j}$ and $w_{j}$ towards the target $\beta$. The total objective on each client is $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{wm}},$ where $\lambda=10^{-4}$ is tuned so that validation accuracy drop stays below 0.1%. Because $\mathcal{L}_{\text{wm}}$ is quadratic and only acts on $|J|/d$ parameters, its contribution to total gradient norm is bounded by $\lambda\beta^{2}p$, guaranteeing negligible drift of the optimisation trajectory.
@@
$S=\frac{1}{|J|}\sum_{j\in J}\theta_{j}\,w_{j},\qquad\text{where }\theta\in\mathbb{R}^{d}\text{ is the released model.}$.
@@
Algorithm Design.
Algorithm 2 gives the client‑side update and verification. Both are compatible with FedAvg and its variants because all clients share $J$ and push updates in the same direction, so every client perturbs the same coordinates with the same sign, when the server averages their updates, those perturbations reinforce one another instead of partially cancelling.
@@
8:     return $\theta$triangleright upload to server
@@
4 Experimental Evaluation and Analysis
This section reports an end‑to‑end implementation of our privacy‑preserving federated learning framework (PPFL‑RDSN) and analyzes how different configuration choices affect efficiency, model quality, and privacy protection. Our evaluation targets two complementary questions:
@@ 4 Experimental Evaluation and Analysis
Each experiment is mapped to one of the above goals and adopts standard super‑resolution metrics (PSNR and SSIM) and exactly the same data splits as prior work.
4.1 Experimental Setup
@@
(I) Simulation platform. All code is written in Python 3.10 with PyTorch 1.12.8 and the CUDA 11.2 toolkit. A single Fedora/Red Hat Linux host emulates the central server and all clients, each client running in a separate thread and bound to its own NVIDIA L40 GPU. Opacus [57] is modified to provide record‑level DP.
@@
(III) Hyper‑parameters. If a client lacks a pretrained model it first initializes $\theta_{0}$ and performs local training. Local RDSN uses 1000 epochs (as in [20]) with Adam, batch size 16, $64^{2}$ LR crops, and an exponentially decaying learning rate starting at $1\times 10^{-4}$. In FL, 10 clients each run $x$ local epochs per round over 10 global rounds, for a total of $10x$ epochs. The overall training time in an ideal environment is:
$T_{\text{total}}=\bigl{(}T_{c_{\text{max}}}+T_{\text{agg}}+T_{\text{updates}}+T_{\text{veri}}\bigr{)}\times T_{\text{glb-rd}},$
where $T_{c_{\text{max}}}$ is the slowest client’s time, $T_{\text{veri}}$ (watermark verification) is negligible for <10 000absent10000<\!10\,000< 10 000 clients, and all other symbols follow Juan et al. [58].
(IV) Baseline timing. Re‑running Wang et al. with two NVIDIA L40 GPUs costs 82,800s for 1000 centralized epochs. Our PPFL‑RDSN converges in only 47 local epochs per round (470 total), cutting wall‑clock time to 77,368s while additionally guaranteeing privacy.
@@
4.2 Performance Evaluation
We adopt the standard super‑resolution benchmarks: Set5, Set14, BSDS100, and Urban100, and the three common scaling factors $\varsigma\in\{2,3,4\}$. Following 470 global epochs on randomly partitioned DIV2K shards, the latest vanilla FL‑RDSN model (no differential privacy (DP) and no watermarking) is broadcast to a dedicated evaluation client that first applies the encryption‑then‑lossy‑compression pipeline before inference. PSNR and SSIM are used throughout as primary quality metrics.
@@ 4.2 Performance Evaluation
Table I compares our vanilla FL‑RDSN with the centralized model of Wang et al.. The two are statistically indistinguishable across all benchmarks and, on larger or more diverse datasets, FL‑RDSN is occasionally better. Per‑image results are visualised in Figures 5 and 6, confirming that the aggregate scores reflect consistent image‑level behavior.
@@
and report PSNR/SSIM in Figure 2. A broad plateau appears at $\epsilon_{\text{best}}\!\in[2.5,3.0]$; we select $\epsilon=2.75$ for all subsequent experiments and tabulate its effect in Table II. Performance remains virtually unchanged, demonstrating that strong formal privacy can be added at negligible cost. Qualitative reconstructions in Section .2 exhibit high perceptual fidelity despite aggressive compression.
@@
(a) $client$ to PSNR, $\varsigma=2$
(b) $client$ to SSIM, $\varsigma=2$
(c) $client$ to PSNR, $\varsigma=3$
(d) $client$ to SSIM, $\varsigma=3$
(e) $client$ to PSNR, $\varsigma=4$
(f) $client$ to SSIM, $\varsigma=4$
@@
(a) $epoch$ to PSNR, $\varsigma=2$
(b) $epoch$ to SSIM, $\varsigma=2$
(c) $epoch$ to PSNR, $\varsigma=3$
(d) $epoch$ to SSIM, $\varsigma=3$
(e) $epoch$ to PSNR, $\varsigma=4$
(f) $epoch$ to SSIM, $\varsigma=4$
@@
5 Related Work
Federated Learning.
Federated Learning (FL) has emerged as a method to train models without centralizing raw data, thereby mitigating concerns about data leakage and expensive data transfers [8]. Unlike early image reconstruction systems that required aggregating data at a single location, FL can handle data spread across multiple clients, each maintaining their own datasets locally. This allows FL-based solutions to cope with diverse image distributions and heterogeneous data sources, which are common in realistic scenarios. However, existing FL approaches for image analysis typically focus on relatively simpler tasks (e.g., classification or segmentation) rather than more complex image reconstruction, where subtle details and high-resolution features are crucial [59, 60, 61]. Moreover, standard FL alone does not ensure that the training process is protected against malicious aggregators or subtle data exposures. For instance, even though FL aims to minimize raw data sharing, model parameters can still reveal private information unless further privacy-enhancing techniques are employed. Also, FL does not inherently protect against model tampering or ensure integrity when updates are aggregated centrally. In contrast, our work designs an FL-based reconstruction framework that not only supports complex, high-fidelity image tasks but also integrates privacy-preserving and model-verification techniques to ensure that neither raw data nor model parameters become an avenue for information leakage or manipulation.
Differential Privacy
Differential Privacy (DP) [9, 10] provides a formal guarantee that the outputs of computations, such as aggregated model updates in Federated Learning (FL), do not reveal sensitive information about any individual input. While applying DP mechanisms to FL has been proposed as a means of mitigating privacy risks in distributed training, standard noise-injection techniques can degrade model accuracy, especially for high-dimensional data such as images. Several works have explored more refined DP mechanisms that adapt noise to internal model states, thereby reducing utility loss [62, 63, 64]. However, these approaches frequently rely on trust assumptions about the execution environment or necessitate complex protocols that introduce overhead, hindering straightforward deployment. Moreover, as with other DP frameworks, vulnerabilities remain: inaccuracies in floating-point arithmetic: [65, 66], precision-based attacks: [67], and misestimation of function sensitivity [68] can compromise DP guarantees. To our knowledge, prior works are not mainly to address the issues in image reconstruction as it will introduce blurriness to the final result, our work deals with this problem smoothly.
Model Integrity and Verification.
As federated learning models become widely shared, ensuring their integrity is crucial. Prior work on model watermarking or fingerprinting focuses on detecting unauthorized copying or identifying the model owner [49]. However, these techniques can also ensure that global models are not replaced or corrupted during training rounds. By embedding markers into the model parameters, clients can verify model authenticity before usage, mitigating the risk of invisible substitutions or malicious alterations.
Extending Encryption-Based Methods.
Encryption-then-lossy-compression approaches have been explored to protect image content from direct inspection [23, 69], but these solutions typically involve only two parties and assume a trusted receiver who trains or applies a reconstruction model. Scaling these approaches to multiple participants introduces new challenges in managing encrypted data and ensuring that all parties trust the global model’s accuracy and authenticity. Achieving both strong data security and effective collaborative model training under these conditions remains non-trivial. The closest to our work is the Residual Dense Spatial Network (RDSN) approach [20], which focuses on image reconstruction quality but does not integrate federated contrastive training [42], enhanced privacy protections (e.i. DP) and verification methods (i.e. Zero-knowledge proofs [70]).
6 Conclusion

### v3

@@ III-A Threat Modeling and Mitigation
  • Local Differential Privacy: Our carefully tailored additive DP noise is applied to high-frequency DCT components of client activations/updates with clipping-based sensitivity control, preventing the inference attacks above by the aggregator or colluding honest-but-curious clients. Per-round $(\varepsilon,\delta)$-LDP composes across rounds via the moments accountant. The detailed design of LDP is described in Section III-B.
  • Model Watermarking Mechanism: As described in Section III-C, a sparse PRF-keyed watermark is embedded during client training and verified by clients before local use of the aggregated model. This prevents model substitution or unauthorized alterations (including replay/rollback) by the aggregator or an active on-path adversary, ensuring that global updates are accepted only if provenance and integrity are verified.
III-B Local Differential Privacy (LDP)
@@
Step 1: Data-independent orthogonal transform.
For a fixed layer $k$, denote its activation
$H=g_{k}(\dots,X;W_{k})\in\mathbb{R}^{C\times h\times w},$
where $g_{k}$ with parameters $W_{k}$ produces a tensor with $C$ channels and spatial size $h\times w$. We apply a data-independent orthogonal transform per channel:
$C_{c}\;=\;TH_{c}T^{\!\top}\quad\text{for }c=1,\dots,C,$
where $H_{c}\in\mathbb{R}^{h\times w}$ is the $c$-th slice and $T\in\mathbb{R}^{h\times h}$ is the (orthonormal) 2-D Discrete Cosine Transform (DCT) matrix along each spatial axis. [^1] Equivalently, $\operatorname{vec}(C_{c})=(T\otimes T)\operatorname{vec}(H_{c})$ with $T\otimes T$ orthonormal. Because $T$ is orthonormal, the transform is an isometry:
$\|C_{c}\|_{F}=\|H_{c}\|_{F}\quad\text{and}\quad\|C_{c}\|_{2}=\|H_{c}\|_{2},$
and thus, for any $\ell_{2}$-based clipping or sensitivity measure that is invariant under orthogonal changes of basis, the sensitivity is unchanged in the transform domain. This norm preservation lets us calibrate DP noise in the DCT domain exactly as in the original domain.
@@
Because the frequency threshold $\tau$ is fixed a priori and independent of private data, the partition is data-independent. By the post-processing invariance of differential privacy, deterministic operations that do not depend on private inputs incur no additional privacy loss [60].
We retain the coefficients in $\mathcal{S}_{L}$ (capturing global structure) and add noise to those in $\mathcal{S}_{H}$ (capturing fine detail). With $C_{L}$ and $C_{H}$ as the masked tensors, we get, $_{L}=T^{\!\top}C_{L}T, HH$$_{H}=T^{\!\top}C_{H}T$.
@@
By the Gaussian mechanism [60] the released $\widetilde{f}$ is $(\epsilon,\delta)$‑DP. The noisy activation is reconstructed as follows:
@@
Algorithm 1: LDP-ClientUpdate.
The full client routine is summarized in Algorithm 1. Multiple rounds are composed via the moments accountant in our experiments, $30$ local steps with $\sigma{=}0.8$ yield a total $\epsilon\!\approx\!3$ at $\delta=10^{-5}$.
@@
[…]
@@
In summary, our client‑side LDP module safeguards the input data while retaining reconstruction quality: we first partition DCT‑domain activations with a publicly fixed threshold, incurring no privacy cost or an explicitly budgeted one if adaptively learned; we then clip the activations and exploit the norm‑preserving orthogonality of the DCT to derive a tight global $L_{2}$ sensitivity of $2C_{\max}$; finally, Gaussian noise calibrated to this bound delivers formal $(\epsilon,\delta)$‑DP guarantees, with privacy loss composed across rounds via the moments accountant. Because noise is confined to high-frequency coefficients that contribute less to overall perceptual fidelity, the dominant low-frequency content remains largely intact, enabling strong privacy protection with minimal degradation in reconstruction quality within the PPFL-RDSN framework.
III-C Model Watermarking
To ensure provenance and integrity of local updates in the PPFL–RDSN framework, we embed a client‑specific digital watermark that (i) does not degrade task performance, (ii) survives federated aggregation, and (iii) remains detectable even in very high‑dimensional models. Our design builds upon two prior approaches: parameter-regularizerr watermarking [68] and double-layer optimization for robustness [69]. Empirically, such schemes incur $0.05\%$ accuracy loss while withstanding pruning of $65\%$ weights or 10 epochs of fine‑tuning, we show detailed results in Table III.
@@ III-C Model Watermarking
Parameter regularization–based watermarking has been extensively studied for centralized neural networks, where a small subset of parameters is regularized toward secret values during training [70, 71]. Such methods, however, rely on a single-owner model and independently chosen watermark locations, making them incompatible with federated learning where parameter updates are aggregated across multiple clients. Recent works on federated watermarking [72] still treat each client’s watermark as isolated, which limits detectability and robustness once the updates are mixed.
To address these limitations, we propose a coordinated sparse model watermark mechanism: all clients share a secret Pseudorandom Function (PRF) key $mk$ to deterministically select an identical, extremely sparse carrier subset $J\subset[d]$, and embed tiny regularization biases only on those coordinates. This design is new in that (i) it preserves client-side privacy while (ii) amplifying watermark detectability at the global-aggregation level, and (iii) maintains indistinguishability from random noise at the single-client level. Below, we formalize how $J$ is sampled and why such coordination remains secure.
@@
where $d$ is the model dimension and $p\!\in\!(10^{-4}-10^{-2})$. We reuse the same secret key $mk$, which makes every client perturb an identical, extremely sparse coordinate subset $J$ chosen by a PRF, so their updates reinforce each other and the watermark emerges clearly only in the aggregated model. This coordination remains secure because the subset $J$ is generated through a secret-keyed pseudorandom function (PRF). Without the key $mk$, an adversary cannot distinguish whether a given index $j$ was intentionally selected or randomly chosen—the distribution of selected indices is computationally indistinguishable from uniform random sampling. Choosing a small sparsity level $p$ limits the number of perturbed coordinates, allowing each client’s per-coordinate watermark amplitude $\alpha$ to be set far below the standard deviation of stochastic gradients $\sigma$. Consequently, the client-side signal-to-noise ratio $\mathrm{SNR}_{\text{client}}=\alpha^{2}/\sigma^{2}$ remains $\ll 1$, meaning that an individual watermark update is statistically indistinguishable from ordinary gradient noise. Because each client’s watermark perturbation is extremely weak ($\mathrm{SNR}_{\text{client}}\!\ll\!1$) and affects only a $p$-fraction of coordinates, identifying the true carrier set $J$ from noisy updates amounts to a high-dimensional sparse-support recovery problem. Hence, $J$ remains effectively hidden even against large collaborative attacks. Thus, sharing the same coordinates is a deliberate design that maximizes watermark detectability while preserving secrecy and robustness.
@@
Instead of post‑hoc noise injection, we jointly optimise the task loss $\mathcal{L}_{\text{task}}$ and a watermark loss $\mathcal{L}_{\text{wm}}=\frac{1}{|J|}\sum_{j\in J}(\theta_{j}w_{j}-\beta)^{2},$ $\text{with }w_{j}\!=\!\text{PRF}(mk,(w,j)),\;\beta>0,$ which drives the correlation between $\theta_{j}$ and $w_{j}$ towards the target $\beta$. The total objective on each client is $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{wm}},$ where $\lambda=10^{-4}$ is tuned so that validation accuracy drop stays below 0.1%. Because $\mathcal{L}_{\text{wm}}$ is quadratic and only acts on $|J|/d$ parameters, its contribution to total gradient norm is bounded by $\lambda\beta^{2}p$, guaranteeing negligible drift of the optimisation trajectory.
@@
$S=\frac{1}{|J|}\sum_{j\in J}\theta_{j}\,w_{j},\qquad\text{where }\theta\in\mathbb{R}^{d}\text{ is the released model.}$
@@
Algorithm 2: PPFL Watermark.
Algorithm 2 outlines both the client-side embedding and the server-side verification procedures. Each client deterministically selects an identical sparse index set $J=\{j\mid\text{PRF}(mk,j)<p\}$ using a shared secret key $mk$, and associates each coordinate $j\!\in\!J$ with a pseudorandom sign $w_{j}=\text{PRF}(mk,("w",j))$. During local training, clients add a small regularization term driving $\theta_{j}w_{j}$ toward a target value $\beta$, ensuring that all clients perturb the same coordinates in the same (signed) direction. Because these updates are aligned across clients, the watermark components reinforce each other when aggregated under FedAvg or its variants, rather than cancel out. In verification, the auditor recomputes $J$ and $w_{j}$ from $mk$, evaluates the normalized correlation score $\frac{1}{|J|}\sum_{j\in J}\theta_{j}w_{j}$, and declares a valid watermark if the score exceeds a detection threshold $\delta$.
@@
8:  return $\theta$ $\triangleright$ upload to server
@@
IV Experimental Evaluation and Analysis
In this section, we present an end‑to‑end implementation of our proposed PPFL‑RDSN, and analyze how different configuration choices affect its efficiency, model quality, and privacy guarantees. Our evaluation targets two complementary questions:
@@ IV Experimental Evaluation and Analysis
Each experiment is mapped to one of the above goals and adopts standard super‑resolution metrics (PSNR and SSIM) and the same data splits as prior works [20, 73, 27].
IV-A Experimental Setup
@@
(I) Simulation platform. All code is written in Python 3.10 with PyTorch 1.12.8 and the CUDA 11.2 toolkit. A single Fedora/Red Hat Linux host emulates the central server and all clients; each client runs in a separate thread and is bound to its own NVIDIA L40 GPU. Opacus [74] is modified to provide record‑level DP.
@@
(III) Hyper‑parameters. If a client lacks a pretrained model, it first initializes $\theta_{0}$ and performs local training. Local RDSN uses 1000 epochs (as in [20]) with Adam optimizer [75], batch size 16, $64^{2}$ LR crops, and an exponentially decaying learning rate starting at $1\times 10^{-4}$. In FL, 10 clients each runs $x$ local epochs per round over 10 global rounds, for a total of $10x$ epochs. The overall training time in an ideal environment is:
$\big[T_{\text{total}}(\varepsilon)\big]\;\approx\;[R_{\varepsilon}]\,\Big([T_{c,\max}]+[T_{\text{agg}}]+[T_{\text{veri}}]\Big)\;+\;T_{\text{init}}\;+\;T_{\text{final}}$
where $T_{c,\max}\;:=\;\max_{i\in S}\Big(T^{\downarrow}_{i}+T^{\text{updates}}_{i}+T^{\uparrow}_{i}\Big)$, which is the slowest client’s time, $T_{\text{veri}}$ (watermark verification) is negligible for <10 000<\!10\,000 clients, and all other symbols follow Juan et al. [76].
(IV) Baseline timing. Re‑running Wang et al. with two NVIDIA L40 GPUs took 82,800s for 1000 centralized epochs. Our PPFL‑RDSN converges in only 47 local epochs per round (470 total), cutting wall‑clock time to 77,368s while additionally guaranteeing privacy.
@@
IV-B Key Management for the Local EtLC Module
For the key sharing in the Local EtLC module, shown in Figure 1, we adopt a standards‑based design that keeps the aggregator untrusted and adds minimal overhead. Each EtLC object is encrypted under a fresh content‑encryption key (CEK) with an Authenticated Encryption with Associated Data (ChaCha20‑Poly1305) [77], and the CEK is delivered to authorized recipients using Hybrid Public Key Encryption (HPKE) [78, 79, 80]. First contact between previously unknown peers can follow Extended Triple Diffie–Hellman to authenticate identity keys and obtain forward secrecy; for 1 to n or dynamic recipients, Messaging Layer Security (MLS) [81] maintains a group secret with forward secrecy and post‑compromise security, from which per‑image CEKs are derived. In scenarios where encrypted EtLC content must later be made accessible to additional clients, proxy re-encryption offers a practical approach: it enables ciphertexts of content-encryption keys to be re-targeted to new recipients by an untrusted proxy, while preserving confidentiality of both the keys and the underlying data. These mechanisms are orthogonal to LDP and watermarking, preserving our training and verification pipeline while providing robust key management.
IV-C Performance Evaluation
We adopt the standard super‑resolution benchmarks: Set5, Set14, BSDS100, and Urban100, and the three common scaling factors $\varsigma\in\{2,3,4\}$. Following 470 global epochs on randomly partitioned DIV2K shards, the latest vanilla FL‑RDSN model (no differential privacy (DP) and no watermarking) is broadcast to a dedicated evaluation client that first applies the encryption‑then‑lossy‑compression pipeline before inference. We use PSNR and SSIM as primary quality metrics. The comparison results are shown in Appendix Sec. -A.
@@ IV-C Performance Evaluation
Table I compares our vanilla FL‑RDSN with the centralized model of Wang et al. They are statistically indistinguishable across all benchmarks. For larger or diverse datasets, FL‑RDSN is occasionally better. Figures 5 and 6 show per‑image results showing that the aggregate scores reflect consistent image‑level behavior.
@@
and report PSNR/SSIM in Figure 2. A broad plateau appears at $\epsilon_{\text{best}}\!\in[2.5,3.0]$; we select $\epsilon=2.75$ for all subsequent experiments and tabulate its effect in Table II. Performance remains virtually unchanged, demonstrating that strong formal privacy can be added at negligible cost. Qualitative reconstructions in Appendix Sec. -B exhibit high perceptual fidelity despite aggressive compression.
@@
(a) clientclient to PSNR, $\varsigma=2$
(b) clientclient to SSIM, $\varsigma=2$
(c) clientclient to PSNR, $\varsigma=3$
(d) clientclient to SSIM, $\varsigma=3$
(e) clientclient to PSNR, $\varsigma=4$
(f) clientclient to SSIM, $\varsigma=4$
@@
(a) epochepoch to PSNR, $\varsigma=2$
(b) epochepoch to SSIM, $\varsigma=2$
(c) epochepoch to PSNR, $\varsigma=3$
(d) epochepoch to SSIM, $\varsigma=3$
(e) epochepoch to PSNR, $\varsigma=4$
(f) epochepoch to SSIM, $\varsigma=4$
@@
V Conclusion and Discussion